OpenStudy (anonymous):

Convert the equation from Rectangular to Polar form: x^2 + y^2 = 9

ganeshie8 (ganeshie8):

use below : \(x^2 + y^2 = r^2\)

ganeshie8 (ganeshie8):

x^2 + y^2 = 9 represents a circle of radius 3, right ?

ganeshie8 (ganeshie8):

u can replace the entire left hand side by r^2 : r^2 = 9 r = 3

ganeshie8 (ganeshie8):

\(r = 3\) is its polar form.